# Boosting HDR Image Reconstruction via Semantic Knowledge Transfer

### v1

@@ Abstract
Recovering High Dynamic Range (HDR) images from multiple Low Dynamic Range (LDR) images becomes challenging when the LDR images exhibit noticeable degradation and missing content. Leveraging scene-specific semantic priors offers a promising solution for restoring heavily degraded regions. However, these priors are typically extracted from sRGB Standard Dynamic Range (SDR) images, the domain/format gap poses a significant challenge when applying it to HDR imaging. To address this issue, we propose a general framework that transfers semantic knowledge derived from SDR domain via self-distillation to boost existing HDR reconstruction. Specifically, the proposed framework first introduces the Semantic Priors Guided Reconstruction Model (SPGRM), which leverages SDR image semantic knowledge to address ill-posed problems in the initial HDR reconstruction results. Subsequently, we leverage a self-distillation mechanism that constrains the color and content information with semantic knowledge, aligning the external outputs between the baseline and SPGRM. Furthermore, to transfer the semantic knowledge of the internal features, we utilize a semantic knowledge alignment module (SKAM) to fill the missing semantic contents with the complementary masks. Extensive experiments demonstrate that our method can significantly improve the HDR imaging quality of existing methods.
1 Intruduction
@@
Multiple-exposure High Dynamic Range (HDR) imaging aims to accurately recover images with a wide luminance range from exposure-varied Low Dynamic Range (LDR) images. However, LDR images are often captured in dynamic scenes, existing deep learning-based methods primarily focus on the aligned LDR sequences and fusing them to obtain high-quality HDR results [31]. Nevertheless, in practical scenarios, LDR images often suffer from noticeable degradation in dark scenes, while bright scenes may experience essential content misses due to occlusion caused by object movement. Despite advancements in recent network designs (e.g., CNNs [12, 37, 49, 25], Transformer-based methods [22, 27, 28, 3]), the inherently ill-posed nature of captured LDR images presents challenges in achieving substantial improvements solely through modifications to network structures.
Figure 1: Performance comparison on the BracketIRE dataset, which consists of RAW format data pairs.
@@
Previous research [32, 45, 34, 18] has demonstrated the potential of using semantic priors to address ill-posed problems in low-level vision tasks. By integrating these priors, models can gain a deeper understanding of image content, providing explicit instructions that guide the reconstruction process. However, no work to date has explored this methodology specifically for multi-exposure HDR imaging. In current low-level vision tasks, the predominant approach [32, 34, 18] involves directly incorporating semantic priors derived from degraded images into the reconstruction network, thereby guiding the model to produce high-quality results. However, these approaches face significant challenges when applied to HDR imaging: First, existing prior extraction models are primarily designed for sRGB-format SDR images. The domain gap between HDR and SDR makes it difficult to yield high-quality priors for HDR imaging, particularly some HDR datasets are stored in RAW format [52]. Second, even when suitable semantic priors are obtained, applying sRGB-derived priors to HDR images in RAW format introduces mismatches due to differences in data representation and dimensionality. Furthermore, HDR imaging relies on multiple LDR images with varying exposures, where semantic information can vary significantly across images, leading to incomplete or conflicting priors from each input.
@@
The aforementioned issues motivate us to leverage semantic priors from another perspective. The motivation for our work is threefold. First, given the variations and conflicts in semantics within each LDR image, we exploit fused HDR images to extract high-quality semantic priors. To this end, we employ a pre-trained semantic segmentation network as a Semantic Knowledge Bank (SKB), which consists of semantic features and segmentation maps. Second, considering the domain gap between HDR and SDR, we perform prior embedding in the SDR domain to enhance the quality of the initial HDR results, fully harnessing the potential of semantic knowledge. Third, inspired by the semantic consistency across different formats and domains, we distill semantic knowledge from the SDR domain to boost the performance of existing HDR models.
@@
Based on the above analysis, we propose a general framework that effectively leverages the advanced capabilities of semantic priors for HDR imaging. Specifically, our proposed framework comprises the Original Reconstruction Model (ORM), the Semantic Priors Guided Reconstruction Model (SPGRM), and the Semantic Knowledge Transfer Scheme (SKTS). The ORM can be any existing HDR imaging method, tasked with generating initial HDR results. The SPGRM utilizes intermediate semantic features derived from the SKB to refine the reconstructed image. Notably, the ORM operates in the HDR domain, whereas the SPGRM functions in the SDR domain. Subsequently, the SKTS exploits semantic consistency across formats/domains, using knowledge distillation [44] to ensure the ORM aligns with the semantic knowledge embodied in the SPGRM. Unlike existing distillation methods that solely distill priors in image content-level, we also introduce a Semantic-Guided Histogram Loss to enhance color consistency across distinct object instances. Furthermore, inspired by MAE’s masked reconstruction [8], we introduce a Semantic Knowledge Alignment Module to assist SKTS, ensuring consistency in the semantic feature space during the distillation process. As illustrated in Fig.1, our framework significantly enhances the performance of all baseline models. The main contributions can be summarized as follows:
@@
  • We propose a general semantic knowledge Transfer framework to boost the performance of existing HDR imaging models without modifying the original network.
  • We propose three key techniques to take full advantage of semantic priors provided by the semantic knowledge bank: the Semantic Priors Guided Reconstruction Model, the Semantic-Guided Histogram Loss, and the Semantic Knowledge Alignment module.
@@
  • We conduct experiments on several RAW/sRGB format datasets, demonstrating that our method can significantly improve the HDR imaging quality of existing methods.
2 Related Work
@@
Multi-exposure HDR Imaging. Traditional methods for multi-exposure HDR imaging primarily rely on preprocessing techniques (e.g., motion rejection, motion registration) [9, 36, 33, 29, 26, 10] to align LDR images and reconstruct high-quality results. However, this approach heavily depends on preprocessing technologies and is less effective for large-scale object movements. In recent years, deep learning-based methods have emerged as a promising alternative, researchers have explored a variety of deep neural network (DNN) architectures and techniques. For instance, attention mechanisms [39, 2, 37] have been employed to selectively focus on relevant image regions, enhancing alignment and fusion accuracy. Similarly, transformer-based models [21, 27, 28, 3] leverage global contextual understanding to improve robustness in HDR reconstruction. Optical flow-based approaches [12, 35, 14] estimate motion between frames to compensate for misalignment, while generative adversarial networks (GANs) [24] and diffusion models [40, 11, 41] have been adapted to synthesize realistic HDR content from imperfect inputs. Additional innovative frameworks, such as [47, 7], further expand the scope of DNN applications in this domain. However, existing multi-exposure HDR imaging methods primarily focus on addressing artifact issues caused by motion, but they overlook the inherently ill-posed nature of captured LDR images in real-world scenarios. When LDR images contain noticeable degradations or information loss, their reconstruction quality is significantly reduced.
@@
Semantic Priors for Image Reconstruction. Recently, methods that utilize semantic priors to guide image reconstruction in low-level visual tasks have proven effective in handling ill-posed problems. Semantic-loss-guided methods [1, 20, 51] employ semantic segmentation loss to explicitly enforce semantic segmentation capabilities within the model. This strategy enhances the model’s capacity to capture deep semantic features. Similarly, Wu et al. [34] and Liang et al. [18] utilize segmentation maps as priors to design tailored loss functions, thereby boosting overall model performance. In contrast, segmentation map [32, 45, 6, 51] and intermediate feature [34, 17, 16] guided methods share similar application strategies, differing only in the type of prior embedded within the model. SFTGAN [32] was one of the earliest methods to use semantic segmentation maps as priors to guide models in low-level visual tasks. Despite their success, existing semantic prior-guided methods are predominantly confined to the SDR domain. This limitation introduces a significant domain gap when extending these techniques to HDR imaging, especially given that HDR images are stored in RAW format. Bridging this gap remains a critical challenge for adapting semantic prior-guided approaches to HDR contexts.
3 Method
3.1 Overview of Proposed Framework
As shown in Fig. 2, our framework primarily consists of the Original Reconstruction Model (ORM) and the Semantic Priors Guided Reconstruction Model (SPGRM), with knowledge self-distillation performed through the Semantic Knowledge Transfer Scheme (SKTS). Given a series of LDR images of a dynamic scene $\left\{L_{1},L_{2},\ldots,L_{n}\right\}$, the ORM utilizes existing HDR reconstruction methods to produce the initial HDR images $H_{in}$. These initial HDR images are then converted into tonemapped sRGB domain results $H_{s}$ through a domain transfer operation:
@@ 3.1 Overview of Proposed Framework
where $\mathcal{T}(\cdot)$ represents the domain transfer operation, which for sRGB format data includes a $\mu$-law function, and for RAW format data additionally includes a differentiable demosaicing operation; $f_{org}$ denotes the ORM network, and $\theta_{o}$ is the network parameters. To extract semantic priors, the subsequent process can be represented as:
@@
where $p$ is the semantic prior, including segmentation maps $p_{m}$ and intermediate features $p_{f}$ with multi-scale dimensions. ${f}_{\text{seg}}$ and $\theta_{s}$ represents the pre-trained segmentation network and its frozen parameters. Then, $p_{f}$ and $H_{s}$ are used as inputs to the SPGRM to obtain the enhanced reconstruction result $\widehat{H_{s}}$:
@@
3.2 Semantic Priors Guided Reconstruction Model
To avoid semantic variations and conflicts in each LDR image, we utilize ORM’s result to extract semantic priors. However, when leveraging these priors to guide the model in reconstructing high-quality images, another challenge that requires careful attention is the disparity between the prior and feature sources. To address this, we implement the SPGRM in the sRGB SDR domain, producing sRGB-tonemapped HDR results. Notably, the ORM and the SPGRM are trained simultaneously, and the transformation operation does not compress any of the original HDR content or impede gradient backpropagation.
As illustrated in Fig. 2, existing HDR imaging models typically follow an “align-then-fuse” paradigm, implicitly aligning LDR images before fusing them to produce HDR results. Given that the SPGRM takes the ORM’s results as input, it solely consists of three Semantic Guided Fusion Modules and excludes the LDR alignment module. In our framework, we employ FastSAM [50] as the SKB owing to its outstanding performance and ease of use. We leverage the output features before the representation head to get four multi-scale semantic features $p_{f}^{i}\in\mathbb{R}^{{C_{i}}\times{\frac{H}{2^{4-i}}}\times{\frac{W}{2^{4-i}}}}(i=0,1,2,3)$. Here, $C_{i}$ represents the channel dimension, and $H$ and $W$ denote the height and width of the input image, respectively. These $p_{f}^{i}$ features are processed through a Feature Pyramid Network (FPN) [19] for multi-level feature interaction, yielding $p_{f}$, which is adjusted to match the channel dimensions of SPGRM’s intermediate features. To incorporate the semantic priors $p_{f}$, we add a Prior Fusion Block to generate the refined feature map $\bar{F_{k}^{T}}$ within each fusion module. Specifically, as depicted in Fig. 2, the Prior Fusion Block takes two inputs: an intermediate feature $F_{k}^{T}$ and a prior feature $p_{f}$. We project $p_{f}$ into a query vector $\mathbf{Q}=W_{d}^{Q}W_{c}^{Q}p_{f}$ using point-wise $1\times 1$ convolution and depth-wise $3\times 3$ convolution with weights $W_{d}^{Q}$ and $W_{c}^{Q}$, respectively. Similarly, $F_{k}^{T}$ is transformed into Key $\mathbf{K}$ and Value $\mathbf{V}$ through analogous operations. Then, we adopt a transposed-attention mechanism [43] to compute the attention map. Hence, the process can be described as follows:

### v2

@@ Abstract
Recovering High Dynamic Range (HDR) images from multiple Standard Dynamic Range (SDR) images become challenging when the SDR images exhibit noticeable degradation and missing content. Leveraging scene-specific semantic priors offers a promising solution for restoring heavily degraded regions. However, these priors are typically extracted from sRGB SDR images, the domain/format gap poses a significant challenge when applying it to HDR imaging. To address this issue, we propose a general framework that transfers semantic knowledge derived from SDR domain via self-distillation to boost existing HDR reconstruction. Specifically, the proposed framework first introduces the Semantic Priors Guided Reconstruction Model (SPGRM), which leverages SDR image semantic knowledge to address ill-posed problems in the initial HDR reconstruction results. Subsequently, we leverage a self-distillation mechanism that constrains the color and content information with semantic knowledge, aligning the external outputs between the baseline and SPGRM. Furthermore, to transfer the semantic knowledge of the internal features, we utilize a Semantic Knowledge Alignment Module (SKAM) to fill the missing semantic contents with the complementary masks. Extensive experiments demonstrate that our framework significantly boosts HDR imaging quality for existing methods without altering the network architecture.
I Introduction
@@
Multiple-exposure High Dynamic Range (HDR) imaging aims to accurately recover images with a wide luminance range from exposure-varied Standard Dynamic Range (SDR) images. However, SDR images are often captured in dynamic scenes, and existing deep-learning-based methods primarily focus on the aligned SDR sequences and fuse them to obtain high-quality HDR results [42]. Nevertheless, in practical scenarios, SDR images often exhibit significant quality degradation in dark scenes due to limited light capture, while bright scenes may suffer from critical content loss caused by motion blur or occlusion, as longer exposure times required for adequate brightness increase susceptibility to object movement. Despite advancements in recent network designs (e.g., CNNs [17, 49, 62, 34], Transformer-based methods [31, 37, 38, 7]), the inherently ill-posed nature of captured SDR images presents challenges in achieving substantial improvements solely through modifications to network structures.
Figure 1: Performance comparison on the BracketIRE dataset, which consists of RAW format data pairs, demonstrates that our framework enables the baseline model to achieve significant improvements in PSNR (pixel accuracy), SSIM (detail preservation), and LPIPS (perceptual fidelity) without altering the network architecture.
@@
Previous research [43, 59, 45, 27] has demonstrated the potential of using semantic priors to address ill-posed problems in low-level vision tasks. By integrating these priors, models can gain a deeper understanding of image content, providing explicit instructions that guide the reconstruction process. However, no work to date has explored this methodology specifically for multi-exposure HDR imaging. In current low-level vision tasks, the predominant approach [43, 45, 27] involves directly incorporating semantic priors derived from degraded images into the reconstruction network, thereby guiding the model to produce high-quality results. However, these approaches face significant challenges when applied to HDR imaging: First, existing prior extraction models are primarily designed for sRGB-format SDR images. The domain gap between HDR and SDR not only complicates the generation of high-quality priors suitable for HDR imaging but also limits their effective integration into the HDR workflow. This challenge is particularly pronounced when handling RAW-format HDR datasets [66], leading to suboptimal performance in HDR tasks. Furthermore, HDR imaging relies on multiple SDR images with varying exposures, where semantic information can vary significantly across images, leading to incomplete or conflicting priors from each input.
@@
The aforementioned issues motivate us to leverage semantic priors from another perspective. The motivation for our work is threefold. First, given the variations and conflicts in semantics within each SDR image, we exploit fused HDR images to extract high-quality semantic priors. To this end, we employ a pre-trained semantic segmentation network as a Semantic Knowledge Bank (SKB), which consists of semantic features and segmentation maps. Second, considering the domain gap between HDR and SDR, we perform prior embedding in the SDR domain to enhance the quality of the initial HDR results, fully harnessing the potential of semantic knowledge. Third, inspired by the semantic consistency across different formats and domains, we distill semantic knowledge from the SDR domain to boost the performance of existing HDR models.
@@
Based on the above analysis, we propose a general framework that effectively leverages the advanced capabilities of semantic priors for HDR imaging. As shown in Fig. 2, our proposed framework comprises the Original Reconstruction Model (ORM), the Semantic Priors Guided Reconstruction Model (SPGRM), and the Semantic Knowledge Transfer Scheme (SKTS). The ORM can be any existing HDR imaging method, tasked with generating initial HDR results. The SPGRM utilizes intermediate semantic features derived from the SKB to refine the reconstructed image. Notably, the ORM operates in the HDR domain, whereas the SPGRM functions in the SDR domain. Subsequently, the SKTS exploits semantic consistency across formats/domains, using knowledge distillation [58] to ensure the ORM aligns with the semantic knowledge embodied in the SPGRM. Unlike existing distillation methods that solely distill priors in the level of image content, we also introduce a Semantic-Guided Histogram Loss to enhance color consistency across distinct object instances. Furthermore, inspired by MAE’s masked reconstruction [13], we introduce a Semantic Knowledge Alignment Module to assist SKTS, ensuring consistency in the semantic feature space during the distillation process. As illustrated in Fig. 1, our framework significantly enhances the performance of all baseline models. The main contributions can be summarized as follows:
@@
  • We propose a general semantic knowledge transfer framework that leverages self-distillation technology to boost the performance of existing multi-exposure HDR reconstruction models, without modifying the original network architecture.
  • We introduce a Semantic-Guided Histogram Loss and a Semantic Knowledge Alignment Module to enable effective semantic knowledge distillation at both the feature and output levels.
@@
  • We conduct experiments on several RAW/sRGB format datasets, demonstrating that our method can significantly improve the HDR imaging quality of existing methods.”
II Related Work
II-A Multi-exposure HDR Imaging
@@
Multi-exposure HDR imaging has been a pivotal area of research in computational photography, aiming to reconstruct high-quality HDR images from a set of SDR images captured at different exposure levels [4]. Traditional methods in this domain have primarily relied on pre-processing techniques such as motion rejection and motion registration [14, 51, 44, 39, 36, 15]. These techniques are designed to align the SDR images and mitigate artifacts caused by motion between exposures. However, their effectiveness diminishes in scenarios involving large-scale object movements or complex scene dynamics, as they heavily depend on the accuracy of preprocessing steps. In recent years, deep learning-based methods have emerged as a promising alternative, researchers have explored a variety of deep neural network (DNN) architectures and techniques. For instance, attention mechanisms [48, 6, 49, 2] have been employed to selectively focus on relevant image regions, enhancing alignment and fusion accuracy. Similarly, transformer-based models [37, 38, 7, 30] leverage global contextual understanding to improve robustness in HDR reconstruction. Optical flow-based approaches [17, 46, 21] estimate motion between frames to compensate for misalignment, while generative adversarial networks (GANs) [35, 22] and diffusion models [16, 52, 50] have been adapted to synthesize realistic HDR content from imperfect inputs. Additional innovative frameworks, such as [57, 11, 25, 33, 47], further expand the scope of DNN applications in this domain. AFUNet[25] adopts a deep-unfolding formulation with cross-iterative coupling between alignment and fusion, which strengthens motion compensation and detail aggregation in dynamic scenes. SMHDR[33] introduces semantic masking together with a curriculum-learning schedule that progressively suppresses saturated or misaligned regions, thereby improving reconstruction robustness. FlexHDR [5] models alignment and exposure uncertainties to produce high quality HDR results. AdaptiveAE[47] proposes an adaptive exposure strategy for dynamic scenes; although orthogonal to network architecture, it complements reconstruction by reducing saturation and ghosting at capture time.
Despite these advancements, existing multi-exposure HDR imaging methods primarily focus on addressing artifacts caused by motion. They often overlook the inherently ill-posed nature of captured SDR images in real-world scenarios, where images may suffer from noticeable degradations or information loss due to factors like noise, blur, or saturation. Consequently, the reconstruction quality of HDR images is significantly reduced when the input SDR images are of suboptimal quality. This limitation underscores the need for methods that can robustly handle degraded inputs while maintaining high reconstruction fidelity.
@@
II-B Semantic Priors for Image Reconstruction.
In the broader context of image reconstruction, particularly in low-level visual tasks such as super-resolution, denoising, and deblurring, the use of semantic priors has emerged as a powerful strategy to address ill-posed problems. Semantic priors provide high-level contextual information that can guide the reconstruction process, enabling models to produce more accurate and perceptually pleasing results. One prominent approach is the use of semantic-loss-guided methods [1, 29, 65, 56], which incorporate semantic segmentation loss into the training process. This strategy explicitly enforces the model’s ability to capture deep semantic features, thereby enhancing its performance in reconstructing images with complex structures. For example, Wu et al.[45] and Liang et al.[27] utilize segmentation maps as priors to design tailored loss functions, which boost the overall performance of the model by aligning the reconstruction with semantic boundaries. Another category of methods involves the direct integration of semantic information into the model’s architecture. Segmentation-map-guided methods [43, 59, 10, 65] and intermediate-feature-guided methods [45, 26, 24] embed semantic priors, such as segmentation maps or intermediate features from pre-trained networks, to condition the reconstruction process. SFTGAN [43], for instance, was one of the earliest methods to use semantic segmentation maps as priors to guide models in low-level visual tasks, demonstrating significant improvements in texture and detail recovery. Semantic priors have also been applied in HDR contexts, such as inverse tone mapping. For example, Goswami et al.[12] use semantic segmentation and an ordered graph to guide diffusion-based inpainting for reconstructing clipped regions in single SDR images, lifting them to HDR.
Despite their success in the SDR domain, these semantic-prior-guided methods face challenges when extended to HDR imaging. A significant limitation is the domain gap between SDR and HDR images, particularly since HDR images are typically stored in RAW format, which differs substantially from the processed SDR images used in most existing methods. This discrepancy necessitates careful adaptation of semantic priors to ensure they are compatible with the unique characteristics of HDR data. Bridging this gap remains a critical challenge for effectively applying semantic prior-guided approaches to HDR contexts.
III Method
III-A Overview of Proposed Framework
As shown in Fig. 2, our framework primarily consists of the Original Reconstruction Model (ORM) and the Semantic Priors Guided Reconstruction Model (SPGRM), with knowledge self-distillation performed through the Semantic Knowledge Transfer Scheme (SKTS). Given a series of SDR images of a dynamic scene $\left\{L_{1},L_{2},\ldots,L_{n}\right\}$, the ORM utilizes existing HDR reconstruction methods to produce the initial HDR images $H_{in}$. These initial HDR images are then converted into tonemapped sRGB domain results $H_{s}$ through a domain transfer operation:
@@ III-A Overview of Proposed Framework
where $\mathcal{T}(\cdot)$ represents the domain transfer operation, which for sRGB format data includes a $\mu$-law function, and for RAW format data additionally includes a differentiable demosaicing operation; $f_{org}$ denotes the ORM network, and $\theta_{o}$ is the network parameter. To extract semantic priors, the subsequent process can be represented as:
@@
where $p$ is the semantic prior, including segmentation maps $p_{m}$ and intermediate features $p_{f}$ with multi-scale dimensions. ${f}_{\text{seg}}$ and $\theta_{s}$ represent the pre-trained segmentation network and its frozen parameters. Then, $p_{f}$ and $H_{s}$ are used as inputs to the SPGRM to obtain the enhanced reconstruction result $\widehat{H_{s}}$:
@@
III-B Semantic Priors Guided Reconstruction Model
To avoid semantic variations and conflicts in each SDR image, we utilize ORM’s result to extract semantic priors. However, when leveraging these priors to guide the model in reconstructing high-quality images, another challenge that requires careful attention is the disparity between the prior and feature sources. To address this, we implement the SPGRM in the sRGB SDR domain, producing sRGB-tonemapped HDR results. Notably, the ORM and the SPGRM are trained simultaneously, and the transformation operation does not compress any of the original HDR content or impede gradient backpropagation.
As illustrated in Fig. 2, existing HDR imaging models typically follow an “align-then-fuse” paradigm, implicitly aligning SDR images before fusing them to produce HDR results. Given that the SPGRM takes the ORM’s results as input, it solely consists of three Semantic Guided Fusion Modules and excludes the SDR alignment module. In our framework, we employ FastSAM [64] as the SKB owing to its outstanding performance and ease of use. Compared to SAM [20], FastSAM achieves inference speeds tens of times faster while maintaining robust semantic prior extraction capabilities. We leverage the output features before the representation head to get four multi-scale semantic features $p_{f}^{i}\in\mathbb{R}^{{C_{i}}\times{\frac{H}{2^{4-i}}}\times{\frac{W}{2^{4-i}}}}(i=0,1,2,3)$. Here, $C_{i}$ represents the channel dimension, and $H$ and $W$ denote the height and width of the input image, respectively. These $p_{f}^{i}$ features are processed through a Feature Pyramid Network (FPN) [28] for multi-level feature interaction, yielding $p_{f}$, which is adjusted to match the channel dimensions of SPGRM’s intermediate features. To incorporate the semantic priors $p_{f}$, we add a Prior Fusion Block to generate the refined feature map $\bar{F_{k}^{T}}$ within each fusion module. Specifically, as depicted in Fig. 2, the Prior Fusion Block takes two inputs: an intermediate feature $F_{k}^{T}$ and a prior feature $p_{f}$. We project $p_{f}$ into a query vector $\mathbf{Q}=W_{d}^{Q}W_{c}^{Q}p_{f}$ using point-wise $1\times 1$ convolution and depth-wise $3\times 3$ convolution with weights $W_{d}^{Q}$ and $W_{c}^{Q}$, respectively. Similarly, $F_{k}^{T}$ is transformed into Key $\mathbf{K}$ and Value $\mathbf{V}$ through analogous operations. Then, we adopt a transposed-attention mechanism [55] to compute the attention map. Hence, the process can be described as follows: